if a 100 kg sample of aluminum at 25.0 °C is heated with 500 kJ of heat energy, what will be the final temperature of the metal sample? The specific heat of aluminum is 0.897 J/ g• °C
Answer #1

q = m S \Delta T

q is the heat supplied (500 kJ or 500,000 J), m is the mass (100 kg or 100,000 g), S is the specific heat (0.897 J/g . C) and \Delta Tis temperature change.

500000 = 100000 \times 0.897 (\Delta T-25) \\ \Delta T = T- 25 = 5.6 \\ T = 30.6 ^oC